# Deblending the MIGHTEE-COSMOS survey with XID+: The resolved radio source counts to Jy

### v1

@@ Abstract
Deep radio continuum surveys provide fundamental constraints on galaxy evolution, but source confusion limits sensitivity to the faintest sources. We present a complete framework for producing high-fidelity deblended radio catalogs from the confused MIGHTEE maps using the probabilistic deblending framework XID+ and prior positions from deep multi-wavelength data in the COSMOS field. To assess performance, we construct MIGHTEE-like simulations based on the Tiered Radio Extragalactic Continuum Simulation (T-RECS) radio source population, ensuring a realistic distribution of star-forming galaxies and active galactic nuclei (AGN) for validation. Through these simulations, we show that prior catalog purity is the dominant factor controlling deblending accuracy: a high-purity prior, containing only sources with a high likelihood of radio detection, recovers accurate flux densities and reproduces input source counts down to (where thermal noise). On the other hand, a complete prior overestimates the source counts due to spurious detections. Our optimal strategy combines the high-purity prior with a mask that removes sources detected above Jy. Applied to the 1.3 deg2 area of the MIGHTEE-COSMOS field defined by overlapping multi-wavelength data, this procedure yields a deblended catalog of 89,562 sources. The derived 1.4 GHz source counts agree with independent P(D) analyses and indicate that we resolve the radio background to Jy. We also define a recommended high-fidelity sample of 20,757 sources, based on detection significance, flux density, and goodness-of-fit, which provides reliable flux densities for individual sources in the confusion-limited regime.
@@
The bright radio sky, at Giga-Hertz frequencies and flux densities above a few millijanskys (mJy), is dominated by a powerful population of radio-loud Active Galactic Nuclei (AGN; e.g., Kellermann1989; Best2012; Hardcastle2020; Whittam2022). In these sources, the observed emission is predominantly non-thermal synchrotron radiation, produced as relativistic electrons, accelerated within powerful jets launched by central supermassive black holes (e.g., BlandfordRees1974; Blandford1979). The statistical distribution of extragalactic sources is an important tool in cosmology, quantified by the differential source counts (), which measure the number of sources per unit solid angle as a function of flux density. At these bright, mJy-to-Jy flux densities, where radio-loud AGN dominate, the normalized counts are relatively flat due to the rarity of these luminous AGN, whose sky density is very low (e.g., Condon1984). However, below a few mJy, they exhibit a well-established upturn, a clear signature that the source population is evolving and was significantly more numerous (e.g., Condon1984; Windhorst1985). This upturn marks a shift in the nature of the dominant radio source population.
@@
The emerging population that dominates the faint sub-mJy and Jy radio sky is now understood to be composed of two main classes of object: star-forming galaxies (SFGs) and radio-faint AGN (e.g., Jarvis2004; Wilman2008; deZotti2010; White2015; White2017; Padovani2016; Smolcic2017). In SFGs, the radio emission at Giga-Hertz frequencies originates from processes linked to massive stars: it is dominated by non-thermal synchrotron radiation from cosmic-ray electrons accelerated in supernova remnants (the endpoints of massive stellar evolution), and a smaller contribution from thermal free-free emission from HII regions, which are clouds of interstellar gas ionized by the intense ultraviolet radiation from young, massive stars (Condon1992; Bell2003; Murphy2011; Kennicutt2012). Because these processes are a direct consequence of recent massive star formation, the radio luminosity serves as a well-established tracer of the star formation rate (SFR, Condon1992). Furthermore, because radio waves are unaffected by the dust that obscures optical and ultraviolet light, these observations provide one of the least biased measures of the cosmic star formation history, an important quantity for galaxy evolution studies (Karim2011; Jarvis2015; Novak2017; Ocran2020; Matthews2021b; Malefahlo2022; Wang2024b; Matthews2024).
@@
The origin of radio emission in the radio-quiet AGN population, is a subject of ongoing debate. Defined as AGN that lack the powerful, large-scale relativistic jets characteristic of their radio-loud counterparts, their radio emission is orders of magnitude weaker. One theory argues that the emission, while weak, is still intrinsically linked to AGN activity, such as small-scale jets, coronal emission from the accretion disk, or outflows and winds (White2015; White2017; Hwang2018; Rankine2021; Wang2025). An alternative theory is that this emission is not directly related to the AGN but is instead dominated by star formation in the host galaxy (Kimball2011; Condon2013; Bonzini2013; Gurkan2015; Malefahlo2020). The current consensus suggests that both processes likely contribute, but their relative importance as a function of key parameters like black hole accretion rate, stellar mass, and redshift remains a key open question in galaxy evolution (e.g. Bonfield2011; Hickox2014; Panessa2019; Macfarlane2021; Yue2024; Kondapally2025).
@@
The radio emission originating from star-formation processes is empirically found to correlate tightly with the infrared luminosity of galaxies, a relationship first discovered in the far-infrared (FIR; (Helou1985; Yun2001) and now commonly referred to as the total infrared–radio correlation (IRRC; e.g., Bell2003). This near-linear correlation holds over several orders of magnitude in luminosity and is known to extend to high redshifts, although its precise evolution with cosmic time remains an active area of research (e.g., Jarvis2010; Magnelli2015; Read2018; Delhaize2017; Delvecchio2021). Given the strong correlation between far-infrared emission and the star-formation rate (SFR), this leads to the radio-SFR relation (e.g. Yun2001; Bell2003; Gurkan2018; Smith2021) and the potential to study star-formation in the highest redshift galaxies (e.g. Whittam2025). There is a growing body of evidence that demonstrates that it is not only pure star-forming galaxies that follow this relation; the host galaxies of most low-luminosity and radio-quiet AGN are also observed to lie on or very near the IRRC (e.g. Moloko2025).
@@
The current generation of radio telescopes has now reached the sensitivity, resolution and survey speed needed to probe the Jy sky in detail. Key instruments include the MeerKAT telescope in South Africa (Jonas2009; Booth2009), the Australian SKA Pathfinder (ASKAP; Johnston2008), and the Low-Frequency Array (LOFAR; vanHaarlem2013). These are enhanced by upgraded arrays like the Giant Metrewave Radio Telescope (uGMRT Gupta2017) , with the future Square Kilometre Array Observatory (SKAO; Braun2015) and the Next Generation Very Large Array (ngVLA; Murphy2018) set to push sensitivity even further. With their ability to combine sub-Jy sensitivity with arcsecond resolution over wide fields, these facilities are designed specifically to address key questions in galaxy evolution: to build a complete census of the cosmic star formation history, disentangle the faint SFG and radio-faint AGN populations, and investigate the faint-end slope of the radio source counts.
@@
Despite these instrumental advances, deep surveys run into the instrumental confusion: the surface density of faint sources becomes so high that individual objects overlap/blend due to the telescope’s beam, making direct detection and photometry unreliable (Condon1974; Zwart2014). There are two strategies that have been developed to push beyond confusion. The first is the probability of deflection or P(D) analysis, which can recover the underlying source counts but does not provide individual source flux densities (Scheuer1957; Condon1974; Condon2012; Vernstrom2014; Matthews2021). The second approach is prior-driven deblending, which uses high-resolution ancillary data to model and partition the confused maps, predominantly for submillimetre and far-infrared data (Tibshirani1996; Zou2006; Magnelli2010; Bethermin2010; Roseboom2010; Chapin2011; Lee2013; Wang2014; Hurley2017; Jin2018; Wang2024). The prior-driven approach has multiple methods, from image-plane model fitting like ‘super-deblending’ (Liu2018; Jin2018) to probabilistic Bayesian frameworks like XID+ (Hurley2017), which samples the full posterior probability distribution of flux densities. However, the success of these probabilistic frameworks is highly sensitive to the construction of the prior lists.
@@
In this paper, we develop and validate a complete framework based on this probabilistic approach to produce a deblended radio source catalog from the confused survey maps of the MeerKAT International GHz Tiered Extragalactic Exploration (MIGHTEE) survey (Jarvis2016; Heywood2022). MIGHTEE is a large survey project using the MeerKAT telescope to map deep extragalactic fields, with the first data release (DR1; Hale2025) providing the data for this work. We use realistic simulations to optimize our methodology before applying this validated strategy to the MIGHTEE observations of the Cosmic Evolution Survey (COSMOS; Scoville2007) field. In Section 2, we describe the MIGHTEE radio data and the ancillary multi-wavelength catalogs used to construct our prior lists. Section 3 provides a brief overview of the XID+ deblending framework. In Section 4, we present the analysis of realistic simulations used to validate and optimize our deblending strategy. In Section 5, we apply this optimized strategy to the MIGHTEE-COSMOS field, presenting the final deblended flux catalog and radio source counts. Finally, we summarize our key findings and conclusions in Section 6. Throughout this paper, we assume a flat CDM cosmology with , , and . Furthermore, we define the radio spectral index () by the power-law relation between flux density () and frequency (), given by .
@@ 2 Data
This section details the primary radio data, the ancillary multi-wavelength catalogs, the star-galaxy classification process, and the final definition of the master galaxy sample used for subsequent analysis.
@@ 2.1 MIGHTEE DR1 Radio Data
We use radio data from the first MIGHTEE data release, DR1 (Hale2025). This release includes deep imaging over 20 deg2 of the COSMOS, XMM-LSS, and CDFS fields. For this work, we focus specifically on the MIGHTEE-COSMOS field, which covers a total area of 4.2 deg2. The DR1 data for this field were produced from 22 individual pointings, totaling 139.6 hours of on-target time at L-band (1.2–1.3 GHz). The final mosaicked images are released at two resolutions: a lower resolution (8.9 arcsec) prioritizing sensitivity and a higher resolution (5.2 arcsec). The thermal noise is 1.6 Jy/beam (lower resolution) and 2.2 Jy/beam (higher resolution), with a median measured total noise across the mosaic of 3.5 Jy/beam (Hale2025). We adopt two distinct noise symbols throughout: for the thermal noise and for the total noise, which includes confusion noise . The depth and area of these maps make them rich with faint sources, leading to source confusion that motivates this work. We apply the deblending technique to the deeper low-resolution data, which suffers more from confusion (Heywood2022).
@@ 2.2 Ancillary Multi-wavelength Data
The probabilistic deblending performed in this work requires an input prior list of potential source positions. We construct our master source list using the deep multi-wavelength catalog presented by Varadaraj2023, which builds upon the data compilation of Adams2021. This catalog is built upon the UltraVISTA DR4 near-infrared (NIR) data (McCracken2012), providing deep Y, J, H, and Ks band photometry. This is complemented by optical coverage from the Hyper Suprime-Cam (HSC) Subaru Strategic Program PDR3 (Aihara2022), as well as data from the Canada-France-Hawaii Telescope (CFHT; Gwyn2012) and Spitzer/IRAC 3.6 and 4.5 m photometry (Fazio2004; Mauduit2012).
@@
To obtain the necessary physical properties for our prior construction, we use the COSMOS2020 photometric catalog (Weaver2022). This is a comprehensive, panchromatic photometric catalog covering the COSMOS field, updated from the COSMOS2015 release (Laigle2016). It includes new, deep imaging data from several key surveys: U-band data from the CLAUDS survey (Sawicki2019), g, r, i, z, y-band data from the HSC Subaru Strategic Program PDR2 (Aihara2019), Y, J, H, -band data from the UltraVISTA DR4 survey (McCracken2012), and updated Spitzer/IRAC data from the Cosmic Dawn Survey (EuclidCollaboration2022). Source detection was performed on a deep, izYJHKs detection image. We use the "Farmer" catalog, which was generated using the Tractor software, where the catalog is limited to the UltraVISTA footprint (Weaver2022).
@@
From this catalog, we use the photometric redshifts and inferred total infrared luminosities ( = rest-frame 8–1000 m luminosity) derived using the EAZY code (Brammer2008). This value is an inferred quantity derived from SED fitting based on the principle of energy balance. Specifically, this calculation is based on the method implemented in the EAZY code, which uses Flexible Stellar Population Synthesis (FSPS) templates (Conroy2009) to model the starlight and estimate the energy absorbed by dust. This absorbed energy is then assumed to be re-radiated in the FIR, following the spectral templates of Magdis2012.
@@
The catalog includes a FLAG_COMBINED column, which identifies sources within a clean area of 1.278 deg2 with full HSC, VISTA, and Spitzer/IRAC coverage and removes sources affected by bright stars or artifacts. We use this flag in the final stage of our analysis (Section 5) when constructing our final prior list. For the prior list construction, we cross-match our master catalog with the full Farmer dataset (964,506 sources) using a search radius of 0.3 arcsec. This results in a matched sample of 650,148 sources, which serves as the basis for our star-galaxy classification.

### v2

@@ Abstract
Deep radio continuum surveys provide fundamental constraints on galaxy evolution, but source confusion limits sensitivity to the faintest sources. We present a complete framework for producing high-fidelity deblended radio catalogues from the confused MIGHTEE maps using the probabilistic deblending framework XID+ and prior positions from deep multi-wavelength data in the COSMOS field. To assess performance, we construct MIGHTEE-like simulations based on the Tiered Radio Extragalactic Continuum Simulation (T-RECS) radio source population, ensuring a realistic distribution of star-forming galaxies and active galactic nuclei (AGN) for validation. Through these simulations, we show that prior catalogue purity is the dominant factor controlling deblending accuracy: a high-purity prior, containing only sources with a high likelihood of radio detection, recovers accurate flux densities and reproduces input source counts down to (where thermal noise). On the other hand, a complete prior overestimates the source counts due to spurious detections. Our optimal strategy combines the high-purity prior with a mask that removes sources detected above Jy. Applied to the 1.3 deg2 area of the MIGHTEE-COSMOS field defined by overlapping multi-wavelength data, this procedure yields a deblended catalogue of 89,562 sources. The derived 1.4 GHz source counts agree with independent P(D) analyses and indicate that we resolve the radio background to Jy. We also define a recommended high-fidelity sample of 20,757 sources, based on detection significance, flux density, and goodness-of-fit, which provides reliable flux densities for individual sources in the confusion-limited regime.
@@
The bright radio sky, at Giga-Hertz frequencies and flux densities above a few millijanskys (mJy), is dominated by a powerful population of radio-loud Active Galactic Nuclei (AGN; e.g., Kellermann et al. 1989; Best and Heckman 2012; Hardcastle and Croston 2020; Whittam et al. 2022). In these sources, the observed emission is predominantly non-thermal synchrotron radiation, produced as relativistic electrons, accelerated within powerful jets launched by central supermassive black holes (e.g., Blandford and Rees 1974; Blandford and Königl 1979). The statistical distribution of extragalactic sources is an important tool in cosmology, quantified by the differential source counts (), which measure the number of sources per unit solid angle as a function of flux density. At these bright, mJy-to-Jy flux densities, where radio-loud AGN dominate, the normalized counts are relatively flat due to the rarity of these luminous AGN, whose sky density is very low (e.g., Condon 1984). However, below a few mJy, they exhibit a well-established upturn, a clear signature that the source population is evolving and was significantly more numerous (e.g., Condon 1984; Windhorst et al. 1985). This upturn marks a shift in the nature of the dominant radio source population.
@@
The emerging population that dominates the faint sub-mJy and Jy radio sky is now understood to be composed of two main classes of object: star-forming galaxies (SFGs) and radio-faint AGN (e.g., Jarvis and Rawlings 2004; Wilman et al. 2008; de Zotti et al. 2010; White et al. 2015, 2017; Padovani 2016; Smolčić et al. 2017). In SFGs, the radio emission at Giga-Hertz frequencies originates from processes linked to massive stars: it is dominated by non-thermal synchrotron radiation from cosmic-ray electrons accelerated in supernova remnants (the endpoints of massive stellar evolution), and a smaller contribution from thermal free-free emission from HII regions, which are clouds of interstellar gas ionized by the intense ultraviolet radiation from young, massive stars (Condon, 1992; Bell, 2003; Murphy et al., 2011; Kennicutt and Evans, 2012). Because these processes are a direct consequence of recent massive star formation, the radio luminosity serves as a well-established tracer of the star formation rate (SFR, Condon 1992). Furthermore, because radio waves are unaffected by the dust that obscures optical and ultraviolet light, these observations provide one of the least biased measures of the cosmic star formation history, an important quantity for galaxy evolution studies (Karim and others, 2011; Jarvis et al., 2015; Novak et al., 2017; Ocran et al., 2020; Matthews et al., 2021a; Malefahlo et al., 2022; Wang et al., 2024b; Matthews et al., 2024).
@@
The origin of radio emission in the radio-quiet AGN population, is a subject of ongoing debate. Defined as AGN that lack the powerful, large-scale relativistic jets characteristic of their radio-loud counterparts, their radio emission is orders of magnitude weaker. One theory argues that the emission, while weak, is still intrinsically linked to AGN activity, such as small-scale jets, coronal emission from the accretion disk, or outflows and winds (White et al., 2015, 2017; Hwang et al., 2018; Rankine et al., 2021; Wang and others, 2025). An alternative theory is that this emission is not directly related to the AGN but is instead dominated by star formation in the host galaxy (Kimball and others, 2011; Condon et al., 2013; Bonzini and others, 2013; Gürkan et al., 2015; Malefahlo et al., 2020). The current consensus suggests that both processes likely contribute, but their relative importance as a function of key parameters like black hole accretion rate, stellar mass, and redshift remains a key open question in galaxy evolution (e.g. Bonfield et al., 2011; Hickox et al., 2014; Panessa et al., 2019; Macfarlane et al., 2021; Yue et al., 2024; Kondapally et al., 2025).
@@
The radio emission originating from star-formation processes is empirically found to correlate tightly with the infrared luminosity of galaxies, a relationship first discovered in the far-infrared (FIR; (Helou et al., 1985; Yun et al., 2001) and now commonly referred to as the total infrared–radio correlation (IRRC; e.g., Bell 2003). This near-linear correlation holds over several orders of magnitude in luminosity and is known to extend to high redshifts, although its precise evolution with cosmic time remains an active area of research (e.g., Jarvis et al. 2010; Magnelli et al. 2015; Read et al. 2018; Delhaize et al. 2017; Delvecchio et al. 2021). Given the strong correlation between far-infrared emission and the star-formation rate (SFR), this leads to the radio-SFR relation (e.g. Yun et al., 2001; Bell, 2003; Gürkan et al., 2018; Smith et al., 2021) and the potential to study star-formation in the highest redshift galaxies (e.g. Whittam et al., 2025). There is a growing body of evidence that demonstrates that it is not only pure star-forming galaxies that follow this relation; the host galaxies of most low-luminosity and radio-quiet AGN are also observed to lie on or very near the IRRC (e.g.,Moloko et al. 2025).
@@
The current generation of radio telescopes has now reached the sensitivity, resolution and survey speed needed to probe the Jy sky in detail. Key instruments include the MeerKAT telescope in South Africa (Jonas, 2009; Booth et al., 2009), the Australian SKA Pathfinder (ASKAP; Johnston and others 2008), and the Low-Frequency Array (LOFAR; van Haarlem and others 2013). These are enhanced by upgraded arrays like the Giant Metrewave Radio Telescope (uGMRT Gupta et al. 2017) , with the future Square Kilometre Array Observatory (SKAO; Braun et al. 2015) and the Next Generation Very Large Array (ngVLA; Murphy and others 2018) set to push sensitivity even further. With their ability to combine sub-Jy sensitivity with arcsecond resolution over wide fields, these facilities are designed specifically to address key questions in galaxy evolution: to build a complete census of the cosmic star formation history, disentangle the faint SFG and radio-faint AGN populations, and investigate the faint-end slope of the radio source counts.
@@
Despite these instrumental advances, deep surveys run into the instrumental confusion: the surface density of faint sources becomes so high that individual objects overlap/blend due to the telescope’s beam, making direct detection and photometry unreliable (Condon, 1974; Zwart et al., 2015). There are two strategies that have been developed to push beyond confusion. The first is the probability of deflection or P(D) analysis, which can recover the underlying source counts but does not provide individual source flux densities (Scheuer, 1957; Condon, 1974; Condon et al., 2012; Vernstrom et al., 2014; Matthews et al., 2021b). The second approach is prior-driven deblending, which uses high-resolution ancillary data to model and partition the confused maps, predominantly for submillimetre and far-infrared data (Tibshirani, 1996; Zou, 2006; Magnelli et al., 2010; Béthermin et al., 2010; Roseboom et al., 2010; Chapin et al., 2011; Lee et al., 2013; Wang et al., 2014; Hurley et al., 2017; Jin et al., 2018; Wang et al., 2024a). The prior-driven approach has multiple methods, from image-plane model fitting like ‘super-deblending’ (Liu et al., 2018; Jin et al., 2018) to probabilistic Bayesian frameworks like XID+ (Hurley et al., 2017), which samples the full posterior probability distribution of flux densities. However, the success of these probabilistic frameworks is highly sensitive to the construction of the prior lists.
@@
In this paper, we develop and validate a complete framework based on this probabilistic approach to produce a deblended radio source catalogue from the confused survey maps of the MeerKAT International GHz Tiered Extragalactic Exploration (MIGHTEE) survey (Jarvis et al., 2017; Heywood et al., 2021). MIGHTEE is a large survey project using the MeerKAT telescope to map deep extragalactic fields, with the first data release (DR1; Hale et al., 2025) providing the data for this work. We use realistic simulations to optimize our methodology before applying this validated strategy to the MIGHTEE observations of the Cosmic Evolution Survey (COSMOS; Scoville et al. 2007) field. In Section 2, we describe the MIGHTEE radio data and the ancillary multi-wavelength catalogues used to construct our prior lists. Section 3 provides a brief overview of the XID+ deblending framework. In Section 4, we present the analysis of realistic simulations used to validate and optimize our deblending strategy. In Section 5, we apply this optimized strategy to the MIGHTEE-COSMOS field, presenting the final deblended flux catalogue and radio source counts. Finally, we summarize our key findings and conclusions in Section 6. Throughout this paper, we assume a flat CDM cosmology with , , and . Furthermore, we define the radio spectral index () by the power-law relation between flux density () and frequency (), given by .
@@ 2 Data
This section details the primary radio data, the ancillary multi-wavelength catalogues, the star-galaxy classification process, and the final definition of the master galaxy sample used for subsequent analysis.
@@ 2.1 MIGHTEE DR1 Radio Data
We use radio data from the first MIGHTEE data release, DR1 (Hale et al., 2025). This release includes deep imaging over 20 deg2 of the COSMOS, XMM-LSS, and CDFS fields. For this work, we focus specifically on the MIGHTEE-COSMOS field, which covers a total area of 4.2 deg2. The DR1 data for this field were produced from 22 individual pointings, totaling 139.6 hours of on-target time at L-band (1.2–1.3 GHz). The final mosaicked images are released at two resolutions: a lower resolution (8.9 arcsec) prioritizing sensitivity and a higher resolution (5.2 arcsec). The thermal noise is 1.6 Jy/beam (lower resolution) and 2.2 Jy/beam (higher resolution), with a median measured total noise across the mosaic of 3.5 Jy/beam (Hale et al., 2025). We adopt two distinct noise symbols throughout: for the thermal noise and for the total noise, which includes confusion noise . The depth and area of these maps make them rich with faint sources, leading to source confusion that motivates this work. We apply the deblending technique to the deeper low-resolution data, which suffers more from confusion (Heywood et al., 2021).
@@ 2.2 Ancillary Multi-wavelength Data
The probabilistic deblending performed in this work requires an input prior list of potential source positions. We construct our master source list using the deep multi-wavelength catalogue presented by Varadaraj et al. (2023), which builds upon the data compilation of Adams et al. (2021). This catalogue is built upon the UltraVISTA DR4 near-infrared (NIR) data (McCracken et al., 2012), providing deep Y, J, H, and Ks band photometry. This is complemented by optical coverage from the Hyper Suprime-Cam (HSC) Subaru Strategic Program PDR3 (Aihara et al., 2022), as well as data from the Canada-France-Hawaii Telescope (CFHT; Gwyn 2012) and Spitzer/IRAC 3.6 and 4.5 m photometry (Fazio et al., 2004; Mauduit et al., 2012).
@@
To obtain the necessary physical properties for our prior construction, we use the COSMOS2020 photometric catalogue (Weaver et al., 2022). This is a comprehensive, panchromatic photometric catalogue covering the COSMOS field, updated from the COSMOS2015 release (Laigle et al., 2016). It includes new, deep imaging data from several key surveys: U-band data from the CLAUDS survey (Sawicki et al., 2019), g, r, i, z, y-band data from the HSC Subaru Strategic Program PDR2 (Aihara et al., 2019), Y, J, H, -band data from the UltraVISTA DR4 survey (McCracken et al., 2012), and updated Spitzer/IRAC data from the Cosmic Dawn Survey (Euclid Collaboration et al., 2022). Source detection was performed on a deep, izYJHKs detection image. We use the FARMER catalogue, which was generated using the Tractor software, where the catalogue is limited to the UltraVISTA footprint (Weaver et al., 2022). The FARMER catalogue is preferred over the CLASSIC catalogue because it uses profile-fitting photometry via The Tractor software, which provides superior deblending in crowded regions compared to the aperture-based photometry used in CLASSIC (Weaver et al., 2022).
@@
From this catalogue, we use the photometric redshifts and inferred total infrared luminosities ( = rest-frame 8–1000 m luminosity) derived using the EAZY code (Brammer et al., 2008). This value is an inferred quantity derived from SED fitting based on the principle of energy balance. Specifically, this calculation is based on the method implemented in the EAZY code, which uses Flexible Stellar Population Synthesis (FSPS) templates (Conroy et al., 2009) to model the starlight and estimate the energy absorbed by dust. This absorbed energy is then assumed to be re-radiated in the FIR, following the spectral templates of Magdis et al. (2012). We note that the COSMOS2020 catalogue relies solely on optical-to-mid-infrared photometry (up to IRAC 4.5 m) and does not include far-infrared or longer wavelength data in the SED fitting process (Weaver et al., 2022).
@@
The catalogue includes a FLAG_COMBINED column, which identifies sources within a clean area of 1.278 deg2 with full HSC, VISTA, and Spitzer/IRAC coverage and removes sources affected by bright stars or artefacts. We use this flag in the final stage of our analysis (Section 5) when constructing our final prior list. For the prior list construction, we cross-match our master catalogue with the full FARMER dataset (964,506 sources) using a search radius of 0.3 arcsec. This results in a matched sample of 650,148 sources, which serves as the basis for our star-galaxy classification.